A cylindrical container with a cross-sectional area of 66.2 cm2 holds a fluid of density 856 kg/m3 . At the bottom of the container the pressure is 116 kPa . Assume Pat = 101 kPa, what is h?
1
Expert's answer
2018-04-09T03:12:15-0400
The hydrostatic pressure
P_b- P_at=gρh
So
h=(P_b- P_at)/gρ=(116×〖10〗^3-101×〖10〗^3)/(9.81×865)=1.77 m